Grade 9General Physics

An object of mass 3kg is kept at a distance of 2m from another object of mass 8kg. At what distance from the object of 3kg, another object of mass 4kg be kept such that the net force of gravitation acting on the object of mass 3kg is zero

Valerie
8 Years agoGrade 9

1 Answer

Saswat Pani
8 Years ago
Force of gravitation is equal to Gm1m2/r^2Hence for F1= F2m1m2/r1^2=m2m3/r2^23×8/2×2=3×4/r2^2r2=√2